OpenStudy (loser66):

The total claim amount for a health insurance policy follows a distribution with density function \(f(x)=(1/1000)e^{-x/1000}\) for x>0. The premium for the policy is set at 100 over the expected total claim amount. If 100 policies are sold, what is the approximate probability that the insurance company will have claims exceeding the premiums collected? Please, help

OpenStudy (dumbcow):

\[EV = \int\limits_0^\infty x f(x) dx\] \[a=\text {total prem} = 100(EV +100)\] \[P(x>a) = \int\limits _a^\infty f(x) dx\]

OpenStudy (loser66):

is it not that a =1100?

OpenStudy (loser66):

Since f is exponential function with parameter \(\alpha =1000\), hence \(E(x) =\alpha =1000\) And the company wants exceed 100, that is 1100, right?

OpenStudy (dumbcow):

yes but 100 policies were sold so multiply that by 100 ---> 110,000 to get total prem collected

OpenStudy (loser66):

yes. I got this part

OpenStudy (loser66):

I am on Central theorem, how to turn it into formula of Central theorem?

OpenStudy (dumbcow):

not sure is the integral for probability using density function not a good enough answer?

OpenStudy (loser66):

I have to use Central theorem to solve this problem to show that I understand the theorem and can apply it to solve problems.

OpenStudy (dumbcow):

you have to standardize the sum of the exponential variables Use the following to get Z-score: \[Z = \frac{S_n - n \mu}{\sigma \sqrt{n}}\] \[\mu = \sigma = 1000\] \[n = 100, S_n = 100*1100\] Find probability using standard normal table \[P(Z>1) \approx 0.16\]